OpenStudy (anonymous):

Find the inverse. Then graph the function and its inverse. f(x) = 3x+3

OpenStudy (anonymous):

The inverse of a function is when the "x" and "y" or "f(x)" and "x" are reversed. y=3x+3 x=3y+3 And then we solve for y!

OpenStudy (anonymous):

i know how to do the first part but i don't know how to solve for y and how to graph

OpenStudy (anonymous):

would it be y= x-3/3

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

y=(x-3)/3, and you would graph it like you would graph anything else.